# Quasi-linear equation $\Delta_{p}v+av^{q}=0$ on manifolds with integral bounded Ricci curvature and geometric applications

### v2

@@ Abstract.
We consider nonexistence and gradient estimate for solutions to $\Delta_{p}v+av^{q}=0$ defined on a complete Riemannian manifold with $\chi$-type Sobolev inequality. A Liouville theorem on this equation is established if the lying manifold $(M,g)$ supports a $\chi$-type Sobolev inequality and the $L^{\frac{\chi}{\chi-1}}$ norm of ${\rm Ric}_{-}(x)$ of $(M,g)$ is bounded from upper by some constant depending on $\dim(M)$, Sobolev constant $\mathbb{S}_{\chi}(M)$ and volume growth order of geodesic ball $B_{r}\subset M$. This extends and improves some conclusions obtained recently by Ciraolo, Farina and Polvara [17], but our method employed in this paper is different from their “P-function” method. In particular, for such manifold with a $\chi$-type Sobolev inequality, we give the lower estimate of volume growth of geodesic ball. If $\chi\leq n/(n-2)$, we also establish the local logarithm gradient estimate for positive solutions to this equation under the condition ${\rm Ric}_{-}(x)$ is $L^{\gamma}$-integrable where $\gamma>{\frac{\chi}{\chi-1}}$.
As topological applications of main results(see Corollory 1.8) we show that for a complete noncompact Riemannian manifold on which the Sobolev inequality (1.8) holds true, $\dim(M)=n\geq 3$ and ${\rm Ric}(x)\geq 0$ outside some geodesic ball $B(o,R_{0})$, there exists a positive constant $C(n)$ depending only on $n$ such that, if
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\mathbb{S}_{\frac{n}{n-2}}(M),$
then $(M,g)$ is of a unique end.
Key words and phrases:
non-linear elliptic equation, gradient estimate, $p$-Laplace
@@ 1. Introduction
$\Delta v+v^{q}=0,$ (1.2)
commonly referred to as the Lane-Emden equation. This equation arises in various branches of mathematics, such as the prescribed scalar curvature problem (for q=(n+2)/(n2)q=(n+2)/(n-2), cf. e.g. [51, 52]), the scalar field equation (cf. [4]), the stationary solutions to Euler’s equation on $\mathbb{S}^{2}$ (cf. [19, 20]) and has been studied extensively in the last half century (cf. e.g. [6, 25, 29, 30, 40, 47, 44, 64]).
@@
$u(x)=\left(a+b\lvert x-x_{0}\rvert^{2}\right)^{-\frac{n-2}{2}},\quad\quad n\left(n-2\right)ab=1,$ (1.3)
where $x_{0}\in\mathbb{R}^{n}$, $a>0$ and $b>0$ (see [14] for the two dimensional case). We refer to [13, 54] and [17] for recent development of classification results for equation (1.2).
@@
$\Delta_{p}v=0,\quad p>1.$ (1.4)
The renowned Cheng-Yau’s logarithm gradient estimate showed that when $p=2$, any solution bounded from above or below to (1.4) is a constant provided the Ricci curvature of the Riemannian manifold is nonnegative (cf. [16]). Kotschwar-Ni [35] proved that any positive $p$-harmonic function on a complete Riemannian manifold with nonnegative sectional curvature is a positive constant. Subsequently, this result was verified to remain correct by Wang and Zhang [59] under the condition of nonnegative Ricci curvature, where the authors applied the Nash-Moser iteration technique and Saloff-Coste’s Sobolev inequality (cf. [50, Theorem 3.1]) to study the gradient estimates of equation (1.4). This gradient estimate was later refined to be sharp by Sung and Wang [56].
@@
$\Delta_{p}u+u^{q}\leq 0.$ (1.6)
Grigor’yan and Sun [30] and Zhang [64] investigated the uniqueness of a nonnegative solution to this inequality for $p=2$. Notably, they utilized the condition of volume growth instead of relying on nonnegative Ricci curvature. Similar results hold for general $p>1$ (see [55]).
@@
$\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv.$ (1.7)
It is well-known that (1.7) holds true for a large class of Cartan-Hadamard manifolds on which the Poincaré inequalities hold, for details we refer to [33]. We also notice that if the $\chi$-type Sobolev inequality holds on $(M,g)$ with $\mathrm{dim}(M)\geq 3$, then $\chi\leq n/(n-2)$ (see Theorem 2.1). Hence, we assume $\chi\in(1,n/(n-2)]$ when $\mathrm{dim}(M)\geq 3$ in the rest of this article. Clearly, (1.7) is just the well-known Sobolev inequality if $\chi={n}/(n-2)$, i.e.
$\mathbb{S}_{\frac{n}{n-2}}(M)\left(\int_{M}f^{\frac{2n}{n-2}}dv\right)^{\frac{n-2}{n}}\leq\int_{M}|\nabla f|^{2}dv,$ (1.8)
which holds true for a large class of complete Riemannian manifolds. Indeed, significant efforts have recently been devoted to studying optimal Sobolev inequalities on Riemannian manifolds (see the survey [27] and its references).
In the sequel, we will use the notation ${\rm Ric}_{-}(\,\cdot\,)$ which is defined as
${\rm Ric}_{-}(x)=\max_{|v|=1,\,v\in T_{x}M}\left\{0,\,-{\rm Ric}_{x}(v,v)\right\},\quad\forall x\in M.$
Recently, Ciraolo, Farina and Polvara (cf. [17]) showed the following conclusions, which can be seen as a generalization of the classical Gidas-Spruck’s result for semilinear equation $\Delta u+u^{q}=0$ defined on Riemannian manifolds with non-negative Ricci curvature to the case of the integral bounded Ricci curvature.
Theorem (Theorem 1.5 in [17]).
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq\frac{1}{48}\left(\frac{n+2}{n-2}-q\right)(q-1)(n-2)\mathbb{S}_{\chi}(M)^{2}.$ (1.9)
@@
${\rm{vol}}(B(o,R))=O\left(R^{2+\frac{8}{q-1}}\right),\ \ \ \text{as}\ \ R\rightarrow\infty.$ (1.10)
@@
It was first observed by the first named author [60, Proposition 2.4], and later by Carron (cf. [12]) and Akutagawa (cf. [1]) independently that if the Sobolev constant of a Riemannian manifold $(M^{n},g)$ is positive, then the volume of the geodesic ball of radius $R$ is larger than $CR^{n}$, where the constant $C$ depends only on $n$ and the Sobolev constant.
The first main result of the present article is that the similar volume growth estimate is also established for Riemannian manifold which enjoys the $\chi$-type Sobolev inequality. Throughout the paper, we shall use the abbreviation $B_{r}$ to denote a geodesic ball of radius $r$ centered at any point on a Riemannian manifold. We now state this result precisely as follows.
@@
It is worthy to point out that $\frac{2\chi}{\chi-1}=n$ if $\chi=\frac{n}{n-2}$ where $n=\dim(M)$. So, we recover the volume estimate obtained in [60]. It is easy to see that there holds
$\frac{2\chi}{\chi-1}>2+\frac{8}{q-1},$
if $q>4\chi-3$. Hence, there is a structural contradiction between the assumption $\chi$-type Sobolev inequality holds” and “volume growth assumption (1.10)” in the above theorem due to Ciraolo-Farina-Polvara. The main contribution of the present paper is that we remove the prior volume growth assumption in this theorem. We now state our result precisely.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that ${\rm{vol}}(B(o,R))=O(R^{\beta^{*}})$ for some $\beta^{*}\geq\frac{2\chi}{\chi-1}>0$ where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,p,q,\beta^{*})$ depending on $n$, $p$, $q$ and $\beta^{*}$ such that, if
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
@@
Roughly speaking, we show that for manifolds which enjoy $\chi$-type Sobolev inequalities, and whose volume growth is strictly less than exponential growth, if the $L^{\chi/(\chi-1)}$ norm of $\rm Ric$$_{-}$ is less than $C^{\prime}\mathbb{S}_{\chi}(M)$, where $C^{\prime}$ depends on the growth order of the volume of geodesic ball, then equation (1.1) does not admit any positive solution.
Moreover, for the case $p=2$ and $a=1$, we conclude the following:
@@
Let (M,g)(M,~g) be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that $\mathrm{vol}\left(B(o,R)\right)=O(R^{\beta^{*}})$ for some $\beta^{*}\geq\frac{2\chi}{\chi-1}>0$ where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,q,\beta^{*})$ depending on $n$, $q$ and $\beta^{*}$ such that, if
@@
In particular, for harmonic or $p$-harmonic functions on a noncompact complete Riemannian manifold enjoying a usually Sobolev inequality we obtain the following direct corollary.
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,p,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
Remark 1.
Here, we want to give some remarks about Theorem 1.2 and Theorem 1.3.
@@
    Our result remove the prior volume growth condition (1.10).
@@
    Compared to Theorem Theorem where the condition $q>1$ is necessary because of the condition (1.9), our results still holds when $q\in(-\infty,1]$.
@@
    We establish a Liouville theorem for the general $p$-Laplace equation (1.1) where $a$ is a general constant.
@@
    In a certain sence, these two theorems can be seen as an effective version of the classical Liouville results due to Gidas-Spruck [29] and Serrin-Zou [53]. To see this, let $g_{\epsilon}$ be a sequence of metrics on $\mathbb{R}^{n}$ with $g_{\epsilon}$ equals to the Euclidean metric outside some compact set of $\mathbb{R}^{n}$ and with $g_{\epsilon}$ converges to the Euclidean metric in $C^{2}$ sense as $\epsilon\rightarrow 0$. Then the Positive Mass Theorem tells us that $g_{\epsilon}$ can not be of nonnegative Ricci curvature. Since the Sobolev inequality only depends on the $C^{0}$ property of the metric. Our results indicate that the classical Liouville property also holds if the deformed metric $g_{\epsilon}$ satisfied
    $\|{\rm Ric}^{g_{\epsilon}}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,p,q)\mathbb{S}_{\frac{n}{n-2}}(\mathbb{R}^{n},g_{\epsilon}).$
It is worth noticing that
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}=\left(\int_{M}{\rm Ric}_{-}^{\frac{n}{2}}(x)dM\right)^{\frac{2}{n}}$
is scale invariant with respect to the metric $g$ equipped on $M$, so the quantity is of obvious and important geometric significance.
If the function ${\rm Ric}_{-}(x)$ belongs to $L^{\gamma}(B_{1})$ for some $\gamma>\chi/(\chi-1)$, by virtue of the relative volume comparison theorem under the integral bounded Ricci curvature condition due to Peterson and Wei (cf. [48]), we obtain the following local gradient estimate for solutions to equation (1.1).
@@
$\sup_{B_{1/2}}\frac{|\nabla v|^{2}}{v^{2}}\leq C\left(p,q,\mathbb{S}_{\chi}(M),\gamma,\|{\rm Ric}_{-}\|_{L^{\gamma}(B_{1})}\right),$ (1.11)
where $\gamma$ is any number greater than $\chi/(\chi-1)$. In particular, if $\|{\rm Ric}_{-}\|_{L^{\gamma}(M)}\leq\Lambda$, then the following global gradient estimate holds
@@
In [49], Petersen and Wei showed that if the Sobolev constant of a manifold is positive and ${\rm Ric}_{-}\in L^{p}$ for some $p>n/2$, then any positive harmonic function $v$ defined on $B_{1}$ satisfies the local gradient estimate
@@
Thus, Theorem 1.5 improves and generalizes Petersen-Wei’s result (cf. Theorem 1.2 in [49]).
Now, we turn to recalling the topological properties of a noncompact complete manifold related to the notion of end. “An end $E$ of a manifold $M$ is an unbounded component of the complement of some compact subset $D$ of $M$. In this case, we say that $E$ is an end corresponding to DD”.
For a noncompact complete manifold with nonnegative Ricci curvature outside a compact set, in 1991 Cai [9] showed that such a manifold is of finitely many ends. Later, in 1995 Cai, Colding and Yang [10] discussed the gap phenomenon for ends of such a class of manifold, concretely, they proved the following theorem:
Given $n>0$, there exists an $\epsilon=\epsilon(n)>0$ such that for all pointed open complete manifolds $(M^{n},o)$ with Ricci curvature bounded from below by $(n-1)\Lambda^{2}$ (for $\Lambda>0$) and nonnegative outside the ball $B(o,a)$, if $\Lambda a<\epsilon(n)$, then $M^{n}$ has at most two ends.
@@
As a geometric application of Corollory 1.4 and Theorem 1.6, we obtain the following gap theorem for ends.
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
It should be pointed out that the conclusion in the above corollary “a complete Riemannian manifold with $\dim(M)=n\geq 3$ and nonnegative Ricci curvature, on which the Sobolev inequality (1.8) holds true, has only an end” is implied by Theorem 3.3 in [60].
@@
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n);$
or more generally, whether or not there exists a positive constant $\epsilon(n)$ depending on $n$ such that any noncompact complete Riemannian manifold is of at most two ends if
$\|{\rm Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n).$
The rest of our paper is organized as follows. In section 2, we will give the volume estimate of a geodesic ball $B_{R}$ in $(M,g)$ on which a $\chi$-type Sobolev inequality holds. Section 3 is devoted to a meticulous estimate of $\mathcal{L}\left(|\nabla\log v|^{2\alpha}\right)$ (the explicit definition of the operator $\mathcal{L}$is given in (3.3)). The proofs of our main results reveal that, by selecting an appropriate parameter $\alpha$, we can establish effective integral estimates for the gradient of positive solutions to equation (1.1). In particular, when $\rm Ric$$_{-}$ satisfies the condition stated in Theorem 1.2 or Theorem 1.5, we obtain a $L^{\theta\chi}$ bound of $|\nabla\log v|$. The crucial thing is this bound only depends on the volume and the radius of the geometric ball. With this integral bound in hand, we then apply the Nash-Moser iteration scheme to complete the proof of Theorem 1.5. In section 4, we continue to prove Theorem 1.3 by choosing suitable auxiliary functions. In Section 5 we provide the proof of Theorem 1.6 and Theorem 1.7.

### v3

@@ Abstract.
We study nonexistence results and gradient estimates for solutions of
$\Delta_{p}v+av^{q}=0$
defined on complete Riemannian manifolds satisfying a $\chi$-type Sobolev inequality. We establish a Liouville theorem under the assumptions that the underlying manifold $(M,g)$ supports a $\chi$-type Sobolev inequality and that the $L^{\frac{\chi}{\chi-1}}$-norm of $\mathrm{Ric}_{-}(x)$ is bounded above by a constant depending only on $\dim(M)$, the Sobolev constant $\mathbb{S}_{\chi}(M)$, and the volume growth rate of geodesic balls $B_{r}\subset M$. This extends and improves several recent results of Ciraolo, Farina, and Polvara [18]; our approach, however, differs from their “$P$-function” method. In addition, for manifolds satisfying a $\chi$-type Sobolev inequality, we obtain a lower bound on the volume growth of geodesic balls. We also derive a local logarithmic gradient estimate for positive solutions, assuming that $\mathrm{Ric}_{-}(x)\in L^{\gamma}$ for some $\gamma>\frac{\chi}{\chi-1}$.
Some geometric and topological applications of our main result are also presented in this article (see Theorem 1.10, Theorem 1.11, and Corollory 1.12). In particular, we prove the following. Let $(M,g)$ be a complete noncompact Riemannian manifold of dimension $n\geq 3$ on which the Sobolev inequality (1.8) holds, and assume that $\mathrm{Ric}(x)\geq 0$ outside some geodesic ball $B(o,R_{0})$. Then there exists a positive constant $C(n)$, depending only on $n$, such that if
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\,\mathbb{S}_{\frac{n}{n-2}}(M),$
then $(M,g)$ has exactly one end.
2010 Mathematics Subject Classification:
Primary 53C21; Secondary 53C20
@@ 1. Introduction
(1.2) $\Delta v+v^{q}=0,$
commonly referred to as the Lane-Emden equation. This equation arises in various branches of mathematics, such as the prescribed scalar curvature problem (for q=(n+2)/(n2)q=(n+2)/(n-2), cf. e.g. [57, 58]), the scalar field equation (cf. [5]), the stationary solutions to Euler’s equation on $\mathbb{S}^{2}$ (cf. [22, 23]) and has been studied extensively in the last half century (cf. e.g. [1, 7, 19, 21, 26, 29, 33, 34, 35, 36, 46, 53, 50, 71]).
@@
(1.3) $u(x)=\left(a+b\lvert x-x_{0}\rvert^{2}\right)^{-\frac{n-2}{2}},\quad\quad n\left(n-2\right)ab=1,$
where $x_{0}\in\mathbb{R}^{n}$, $a>0$ and $b>0$ (cf. [15] for the two dimensional case). We refer to [19, 60] and [18] for recent development of classification results for equation (1.2).
@@
(1.4) $\Delta_{p}v=0,\quad p>1.$
The renowned Cheng-Yau’s logarithm gradient estimate showed that when $p=2$, any solution bounded from above or below to (1.4) is a constant provided the Ricci curvature of the Riemannian manifold is nonnegative (cf. [17]). Kotschwar-Ni [41] proved that any positive $p$-harmonic function on a complete Riemannian manifold with nonnegative sectional curvature is a positive constant. Subsequently, this result was verified to remain correct by Wang and Zhang [65] under the condition of nonnegative Ricci curvature, where the authors applied the Nash-Moser iteration technique and Saloff-Coste’s Sobolev inequality (cf. [56, Theorem 3.1]) to study the gradient estimates of equation (1.4). This gradient estimate was later refined to be sharp by Sung and Wang [62].
@@
(1.6) $\Delta_{p}u+u^{q}\leq 0.$
Grigor’yan and Sun [36] and Zhang [71] investigated the uniqueness of a nonnegative solution to this inequality for $p=2$. Notably, they utilized the condition of volume growth instead of relying on nonnegative Ricci curvature. Similar results hold for general $p>1$ (cf. [61]).
@@
(1.7) $\mathbb{S}_{\chi}(M)\left(\int_{M}f^{2\chi}dv\right)^{\frac{1}{\chi}}\leq\int_{M}|\nabla f|^{2}dv.$
It is well known that (1.7) holds for a large class of Cartan–Hadamard manifolds on which Poincaré inequalities are valid; for details, we refer to [39]. We also note that if a $\chi$-type Sobolev inequality holds on $(M,g)$ with $\dim(M)=n\geq 3$, then necessarily $\chi\leq\frac{n}{n-2}$ (see Theorem 2.1). Therefore, throughout the remainder of this article we assume $\chi\in\bigl(1,\frac{n}{n-2}\bigr]$ when $\dim(M)\geq 3$.
Clearly, (1.7) reduces to the classical Sobolev inequality when $\chi=\frac{n}{n-2}$, namely,
(1.8) $\mathbb{S}_{\frac{n}{n-2}}(M)\left(\int_{M}f^{\frac{2n}{n-2}}\,dv\right)^{\frac{n-2}{n}}\leq\int_{M}|\nabla f|^{2}\,dv,$
which holds for a broad class of complete Riemannian manifolds. In recent years, significant effort has been devoted to optimal Sobolev inequalities on Riemannian manifolds; see, for instance, the survey [31] and the references therein.
In the sequel, we will use the notation $\mathrm{Ric}_{-}(\,\cdot\,)$ which is defined as
$\mathrm{Ric}_{-}(x)=\max_{|v|=1,\,v\in T_{x}M}\left\{0,\,-\mathrm{Ric}_{x}(v,v)\right\},\quad\forall x\in M.$
Recently, Ciraolo, Farina, and Polvara (cf. [18]) proved the following results, which may be viewed as a generalization of the classical Gidas–Spruck’s theorem for the semilinear equation $\Delta u+u^{q}=0$ on Riemannian manifolds with nonnegative Ricci curvature to the setting where the Ricci curvature is only bounded in an integral sense.
Theorem 1.2 (Theorem 1.5 in [18]).
@@
(1.9) $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq\frac{1}{48}\left(\frac{n+2}{n-2}-q\right)(q-1)(n-2)\mathbb{S}_{\chi}(M)^{2}.$
@@
(1.10) ${\rm{Vol}}(B(o,R))=O\left(R^{2+\frac{8}{q-1}}\right),\ \ \ \text{as}\ \ R\rightarrow\infty.$
@@
It was first observed by the first-named author [67, Proposition 2.4], and later independently by Carron (cf. [13]) and Akutagawa (cf. [2]), that if a Riemannian manifold $(M^{n},g)$ has positive Sobolev constant, then the volume of any geodesic ball of radius $R$ satisfies
${\rm{Vol}}(B_{R})\geq CR^{n},$
where the constant $C$ depends only on $n$ and the Sobolev constant.
The first main result of the present article shows that an analogous volume growth estimate also holds for Riemannian manifolds satisfying the $\chi$-type Sobolev inequality. Throughout the paper, we use the abbreviation $B_{r}$ to denote a geodesic ball of radius $r$ centered at an arbitrary point of the manifold. We now state the result precisely as follows.
@@
It is worth pointing out that $\frac{2\chi}{\chi-1}=n$ when $\chi=\frac{n}{n-2}$, where $n=\dim(M)$. In this case, we recover the volume estimate obtained in [67]. Moreover, it is easy to see that
$\frac{2\chi}{\chi-1}>2+\frac{8}{q-1}$
whenever $q>4\chi-3$. Consequently, for such $q$ there is a structural incompatibility between the assumptions that a $\chi$-type Sobolev inequality holds and that the volume growth condition (1.10) is satisfied in the theorem of Ciraolo–Farina–Polvara. The main contribution of the present paper is to remove the a priori volume growth assumption from that theorem. We now state our result precisely.
@@
Let $(M,g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that ${\rm{Vol}}(B(o,R))=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ in the case $\chi=\frac{n}{n-2}$, where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,p,q,\beta^{*})$ depending on $n$, $p$, $q$ and $\beta^{*}$ such that, if
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}\leq C(n,p,q,\beta^{*})\mathbb{S}_{\chi}(M),$
@@
Roughly speaking, we show that if $(M,g)$ satisfies a $\chi$-type Sobolev inequality and has volume growth strictly subexponential, then, provided the $L^{\chi/(\chi-1)}$-norm of $\mathrm{Ric}_{-}$ is bounded by $C^{\prime}\,\mathbb{S}_{\chi}(M)$, where $C^{\prime}$ depends on the growth rate of the volume of geodesic balls, equation (1.1) admits no positive solutions.
Moreover, in the case $p=2$ and $a=1$, we obtain the following conclusion:
@@
Let $(M,\penalty 10000\ g)$ be a complete noncompact Riemannian manifold on which the $\chi$-type Sobolev inequality holds. Assume that $\mathrm{Vol}\left(B(o,R)\right)=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$ in the case $\chi=\frac{n}{n-2}$, where $B(o,R)\subset M$ is a geodesic ball centered at a fixed point $o\in M$. Then there exists a positive constant $C(n,q,\beta^{*})$ depending on $n$, $q$ and $\beta^{*}$ such that, if
@@
Remark 1.6.
Note that $\frac{\chi}{\chi-1}>\frac{n}{2}$ whenever $\chi\in\bigl(1,\,\frac{n}{n-2}\bigr)$. By the relative volume comparison theorem under integral Ricci curvature bounds due to Petersen and Wei [54], if $\|\mathrm{Ric}_{-}\|_{L^{\frac{\chi}{\chi-1}}}$ is bounded, then
$\mathrm{Vol}\bigl(B(o,R)\bigr)=O\!\left(R^{\frac{2\chi}{\chi-1}}\right)\quad\text{as}\quad R\rightarrow\infty.$
Consequently, the assumption “ $\mathrm{Vol}\bigl(B(o,R)\bigr)=O(R^{\beta^{*}})$ for some $\beta^{*}\geq n$” in Theorem 1.4 and Theorem 1.5 is needed only in the borderline case $\chi=\frac{n}{n-2}$.
As a direct corollary of Theorem 1.4, we obtain the following Liouville type result for $p$ harmonic functions.
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,p,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
Remark 1.8.
Here, we record several remarks on Theorem 1.4 and Theorem 1.5.
@@
    Our results remove the prior volume growth assumption (1.10).
@@
    In Theorem 1.2, the restriction $q>1$ is required due to the curvature assumption (1.9). In contrast, our results remain valid for all $q\in(-\infty,1]$.
@@
    We establish a Liouville theorem for the general $p$-Laplace equation (1.1), where $a$ is an arbitrary constant.
@@
    In a certain sense, these two theorems may be viewed as effective versions of the classical Liouville results of Gidas–Spruck [35] and Serrin–Zou [59]. To explain this, let $g_{\varepsilon}$ be a family of metrics on $\mathbb{R}^{n}$ such that $g_{\varepsilon}$ agrees with the Euclidean metric outside a compact set and $g_{\varepsilon}\to g_{\mathrm{Euc}}$ in $C^{2}$ as $\varepsilon\to 0$. The Positive Mass Theorem implies that $g_{\varepsilon}$ cannot have nonnegative Ricci curvature. Since the Sobolev constant depends only on the $C^{0}$-geometry of the metric, our results show that the classical Liouville property still holds for such a deformation provided that
    $\bigl\|\mathrm{Ric}^{g_{\varepsilon}}_{-}\bigr\|_{L^{\frac{n}{2}}}\leq C(n,p,q)\,\mathbb{S}_{\frac{n}{n-2}}(\mathbb{R}^{n},g_{\varepsilon}).$
Assume that $\mathrm{Ric}_{-}(x)\in L^{\gamma}(B_{1})$ for some $\gamma>\frac{\chi}{\chi-1}$. By the relative volume comparison theorem under integral Ricci curvature bounds due to Petersen and Wei (cf. [54]), we obtain the following local gradient estimate for solutions of (1.1).
@@
(1.11) $\sup_{B_{1/2}}\frac{|\nabla v|^{2}}{v^{2}}\leq C\left(p,q,\mathbb{S}_{\chi}(M),\gamma,\|\mathrm{Ric}_{-}\|_{L^{\gamma}(B_{1})}\right),$
where $\gamma$ is any number greater than $\chi/(\chi-1)$. In particular, if $\|\mathrm{Ric}_{-}\|_{L^{\gamma}(M)}\leq\Lambda$, then the following global gradient estimate holds
@@
In [55, Theorem 1.2 ], Petersen and Wei proved that if a manifold has positive Sobolev constant and $\mathrm{Ric}_{-}\in L^{p}$ for some $p>\frac{n}{2}$, then any positive harmonic function $v$ on $B_{1}$ satisfies the local gradient estimate
@@
Therefore, Theorem 1.9 both improves and extends the result of Petersen and Wei.
Now, we turn to recalling some topological features of noncompact complete manifolds, with a focus on estimates for the number of ends. Notice that “an end $E$ of a manifold $M$ is an unbounded component of the complement of some compact subset $D$ of $M$. In this case, we say that $E$ is an end corresponding to DD”.
For a noncompact complete manifold whose Ricci curvature is nonnegative outside a compact set, Cai [10] proved that such a manifold has only finitely many ends. Later, in 1995, Cai, Colding, and Yang [11] studied a gap phenomenon for the ends of this class of manifolds. More precisely, they proved the following theorem:
Given $n>0$, there exists an $\epsilon(n)>0$ such that for all pointed open complete manifolds $(M^{n},o)$ with Ricci curvature bounded from below by $(n-1)\Lambda^{2}$ and nonnegative outside the ball $B(o,a)$, if $\Lambda a<\epsilon(n)$, then $M^{n}$ has at most two ends.
@@
As a geometric application of Corollory 1.7 and Theorem 1.10, we obtain the following gap theorem for ends.
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n,\beta^{*})\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq C(n)\mathbb{S}_{\frac{n}{n-2}}(M),$
@@
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n);$
or more generally, does there exist $\epsilon(n)>0$ such that any noncompact complete Riemannian manifold has at most two ends whenever
$\|\mathrm{Ric}_{-}\|_{L^{\frac{n}{2}}}\leq\epsilon(n)?$
The rest of the paper is organized as follows. In Section 2, we derive a volume estimate for geodesic balls $B_{R}$ in $(M,g)$ under the assumption that $(M,g)$ satisfies a $\chi$-type Sobolev inequality. Section 3 is devoted to a careful estimate of ${\mathscr{L}}\!\left(|\nabla\log v|^{2\alpha}\right)$ (the operator $\mathscr{L}$} is defined explicitly in (3.3)). Our proofs show that, by choosing a suitable parameter $\alpha$, one can obtain effective integral estimates for the gradient of positive solutions to (1.1). In particular, under the assumptions of Theorem 1.4 or Theorem 1.9, we obtain an $L^{\theta\chi}$ bound for $|\nabla\log v|$. Crucially, this bound depends only on the radius and the volume of the geodesic ball. With this integral bound in hand, we then apply the Nash–Moser iteration scheme to complete the proof of Theorem 1.9. In Section 4, we proceed to prove Theorem 1.5 by choosing appropriate auxiliary functions. Finally, in Section 5 we prove Theorem 1.10 and Theorem 1.11.